Pre Algeba- Unit Conversion Application Problems.
The dimersions of Drew and Jills living room are 6 yards by 5 yards. How many square feet of carpet do they need to buy? bIf they are going to re-carpet their living room themselves and the carpet copst $2.95/square foot plus an addition cost of $1.00/square foot for the padding, what will be the approximate total cost of the carpeting (with padding including)before tax?

Answer by mananth(16946)

living room size= 6 by 5 yards
= 18 by 15 feet
area in square feet = 270
cost of carpet + padding = 3.95 / sq. ft
so total cost = 270* 3.95 = $1066.50